OpenStudy (anonymous):

What is the value of p? x=-1/8(y^2) i don't understand this because there is no p in the equation

OpenStudy (anonymous):

it can be peak. if it is, answer is (0,0). but i'm not sure :)

OpenStudy (anonymous):

p? stands for

OpenStudy (anonymous):

the vertex is at (0,0) i got this does this make sense to you? i dont know if my work is right a=-1/8, 1/4p -1/8=1/4p -4p=8 p=-2

OpenStudy (anonymous):

i'm not sure what it stands for.. it doesnt say

OpenStudy (anonymous):

no! why 1/4p=-1/8?!?

OpenStudy (anonymous):

i found that solution online

OpenStudy (anonymous):

\[y ^{2}=-8x\] \[y ^{^{2}}=4(-2)x\] \[y ^{2}=4ax\] a=-2

OpenStudy (anonymous):

of if u treat it as \[y ^{2}=-4ax\] then a=2

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

it can be true. i don't know.

OpenStudy (anonymous):

thanks :) can one of you please help me find the length of the focal width for the same equation? x=-1/8(y^2)

OpenStudy (anonymous):

@marsss @nitz

OpenStudy (anonymous):

ya can u please wait a sec

OpenStudy (anonymous):

you should plug x=-2 and x=2 in equaiton and then find the difference

OpenStudy (anonymous):

sure thanks :) @nitz

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

ill try plugging in x one minute

OpenStudy (anonymous):

when x=2 y=4i,-4i

OpenStudy (anonymous):

when x=y=4,-4

OpenStudy (anonymous):

@marsss what is the diff

OpenStudy (anonymous):

difference*

OpenStudy (anonymous):

length is 4a

OpenStudy (anonymous):

i'm sorry i had a mistake. x does'nt equal 2. when x=2, difference is 4-(-4)=8 ok??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey a and p are one and the same thing....just notational difference

OpenStudy (anonymous):

oh ok, so what would the fi

OpenStudy (anonymous):

final answer be in this case

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

i dont know how to find it from here

OpenStudy (anonymous):

length is 8

OpenStudy (anonymous):

thank you so much :)